A proton travels at 1.2x10^7 m/s in a plane perpendicular to a uniform 0.020T magnetic field. Give a quantitativ description of its path


1
Expert's answer
2021-05-24T14:14:41-0400

Gives

m="1.67\times10^{-27}kg"

"V=1.2\times10^7m/sec"

Magnetic field (B)=0.020T

"\theta=90°"

We know that

Circuler path traced by the proton

"\frac{mv^2}{r}=q(v\times B)"


"\frac{mv^2}{r}=qvBsin90°"

"\frac{mv^2}{r}=qvB"

"r=\frac{mv}{qB}"

Put value


"r=\frac{1.67\times10^{-27}\times{1.2\times10^7}}{{1.6\times10^{-19}\times0.02}}=6.2625m"


Radius of circuler path (r)=6.2625 m